# Is In-Context Learning a Type of Gradient-Based Learning? Evidence from the Inverse Frequency Effect in Structural Priming

### v1

@@ Abstract
Large language models (LLMs) have shown the emergent capability of in-context learning (ICL). One line of research has explained ICL as functionally performing gradient descent. In this paper, we introduce a new way of diagnosing whether ICL is functionally equivalent to gradient-based learning. Our approach is based on the inverse frequency effect (IFE)—a phenomenon in which an error-driven learner is expected to show larger updates when trained on infrequent examples than frequent ones. The IFE has previously been studied in psycholinguistics because humans show this effect in the context of structural priming (the tendency for people to produce sentence structures they have encountered recently); the IFE has been used as evidence that human structural priming must involve error-driven learning mechanisms. In our experiments, we simulated structural priming within ICL and found that LLMs display the IFE, with the effect being stronger in larger models. We conclude that ICL is indeed a type of gradient-based learning, supporting the hypothesis that a gradient component is implicitly computed in the forward pass during ICL. Our results suggest that both humans and LLMs make use of gradient-based, error-driven processing mechanisms.
Is In-Context Learning a Type of Gradient-Based Learning? Evidence from the Inverse Frequency Effect in Structural Priming
@@ 1 Introduction
To what extent do humans and language models use similar processing mechanisms? This question is of interest to both Artificial Intelligence researchers and cognitive scientists. Language models and human learners have some substantial differences: human learners often display a flexible learning ability to adapt to new examples, while language models typically require massive amounts of training data and a large number of parameters to exhibit human-like performance. Yet recent pre-trained large language models (LLMs) have shown the capacity for in-context learning (ICL): they adapt to specific tasks with a few demonstration-answer pairs provided as prompts in the context window without any parameter updates (Brown et al., 2020). This intriguing, emergent capability could provide a way to bridge the divide between language models and human learners: perhaps ICL is a processing mechanism that, like humans, can flexibly adapt to new demands.
@@
Among the body of research into the sources and interpretations of the ICL capability, one line of work has aimed to deepen the theoretical understanding of ICL by offering functional interpretations of ICL as a kind of implicit gradient descent during inference. Garg et al. (2022), Zhang et al. (2024), and Ahn et al. (2024) have shown that standard Transformers (Vaswani et al., 2017) can be trained to implement learning algorithms for linear regressions under ICL-based training objectives. Von Oswald et al. (2023) have demonstrated that Transformer models, with appropriate choices of parameters, can process in-context demonstrations in a way that is functionally equivalent to performing gradient updates on the same demonstration examples. Dai et al. (2023) provided a mathematical construction showing the dual form between Transformer attention and gradient descent and interpreted ICL as a meta-optimization process that performs implicit fine-tuning. However, Shen et al. (2023) observe that the importance of these theoretical demonstrations is limited in that they do not take ICL to be an emergent property, instead assuming a training objective that optimizes for ICL. They therefore deviate from actual LLMs pre-trained with natural data with a language modeling training objective. Indeed, Shen et al. found inconsistencies between ICL and gradient descent in real models, and left the equivalence between ICL and gradient descent an open question.
@@
In this paper, we aim to better characterize what kind of learning mechanism ICL is by drawing a connection between ICL and human learning mechanisms. Specifically, we examine the hypothesis that ICL functionally performs gradient-based fine-tuning (e.g., gradient descent) by empirically investigating a weaker claim with off-the-shelf LLMs evaluated with natural language data: whether ICL is a type of gradient-based—i.e., error-driven—learning such that a gradient component is implicitly computed during the forward pass. We approach this question by treating ICL as a processing mechanism of LLMs and borrowing insights from methods of studying processing mechanisms in humans: we examine to what extent LLMs show the inverse frequency effect (IFE), a phenomenon in the human structural priming paradigm (Branigan and Pickering, 2017) that has been argued to require a gradient-based processing mechanism in humans, namely implicit learning (e.g., Chang et al., 2006). We study the linguistic phenomenon of the dative alternation and demonstrate that LLMs show a robust IFE under standard fine-tuning (which involves explicit gradient-based learning), and varying degrees of the IFE under the ICL setting, with larger models showing a stronger IFE. We conclude that, at least in the cases we studied, ICL is indeed a gradient-based learning mechanism.
@@
Our study has implications for both NLP/machine learning (1 and 1) and linguistically-motivated analysis of LLMs (1 and 1):
@@
. We find evidence that ICL in off-the-shelf LLMs can be viewed as a form of gradient-based learning.
@@
. We show that LLMs qualitatively display an important property of human language processing, namely the IFE in structural priming.
@@
Overall, our results suggest that error-driven learning is an aspect of processing that is shared between humans and LLMs.
@@ 2 Background and Related Work
In this section, we lay out the building blocks necessary for motivating why we use the IFE to diagnose the gradient-based nature of ICL. Our experimental approach is formally described in Section 3.1.
Figure 1: Reasoning behind our current study. Abbreviations: LLMs = large language models; ICL = in-context learning; IFE = the inverse frequency effect (the tendency for error-driven learners to show larger updates when trained on a low-frequency example than a high-frequency one; see text).
@@ 2.1 Structural Priming in Psycholinguistics
Structural priming refers to the phenomenon that speakers tend to reuse recently encountered syntactic structures Bock (1986). For example, after encountering a double object (DO) sentence (e.g., Alice gave Bob a book), speakers tend to produce another DO structure (e.g., The student sent the professor a letter) rather than a prepositional dative (PD) structure (e.g., The student sent a letter to the professor). Structural priming has been interpreted as an adaptation mechanism, where speakers adapt lexical and syntactic predictions to the current context Jaeger and Snider (2013), similar to the way in which LLMs adapt their outputs on the basis of prompts. Examples of priming typically involve two sentences; we will refer to them as the prime sentence and the target sentence, where the prime sentence comes first and influences predictions about the target sentence.
@@
One important aspect of structural priming is the inverse frequency effect Jaeger and Snider (2008); Bernolet and Hartsuiker (2010); Kaschak et al. (2011): less preferred syntactic alternatives (measured by the relative frequency in the speaker’s experience against their counterparts) cause stronger overall priming than more preferred structures. This can be seen by looking at sentences involving verbs with gradient structural preferences, or verb biases (or alternation biases; see Hawkins et al., 2020 for a systematic investigation of verb biases in neural models). The verb give allows both prepositional dative (PD) (The doctor gave a book to the judge) and double object (DO) (The doctor gave the judge a book) structures, but is biased toward (i.e., occurs more often with) DO in English. Under the IFE, a “prime” sentence involving give in its (less preferred) PD structure will cause a greater priming effect than a DO sentence, i.e., a PD prime will increase the probability of a PD occurring more than a DO prime will increase the probability of a DO occuring. The strength of PD priming (i.e., the increase in the probability of a PD target given a PD prime) inversely correlates with the expectation of a PD prime, determined by its verb bias Bernolet and Hartsuiker (2010).
@@
Two mainstream theories have been proposed to account for structural priming. Transient activation theory Pickering and Branigan (1998) claims that the activation of structural representations from the prime persists for a short time (in working memory), such that the structural information has a higher probability of being reactivated on the next relevant opportunity. The current form of transient activation theory does not account for the IFE because the amount of residual activation is independent from verb biases and does not involve any error-driven mechanism. In contrast, implicit learning theory Chang et al. (2006) claims that humans implicitly learn and dynamically adjust probabilistic information about different structures (including verb biases) from experience (in long-term memory) and use such information to predict the form of prime sentences. Crucially, under standard theories of learning, the update performed by the learner is error-driven, such that a larger update is performed in situations where the learner’s predictions are farther from the truth. In the context of priming, this would mean that priming strength is determined by the difference between the learner’s predictions and the actual prime sentence: the less the learner expects the observed prime structure, the larger the gradient is, resulting in a larger priming strength. Therefore, the implicit learning theory—unlike transient activation—predicts the IFE. The two theories are not mutually exclusive and can co-exist to account for priming, a framing known as the dual mechanism account Tooley and Traxler (2010).
@@
In this study, we assume the correctness of the arguments from psycholinguistics that state that only error-driven learning mechanisms could predict the IFE. Therefore, by examining whether LLMs show the IFE in the ICL setting, we can infer whether some type of gradient component is computed in the forward pass without explicit weight updates, which informs us about whether there is an implicit gradient-based component in ICL.
2.2 Connections among Distributional Properties of Pre-Training Data, Priming, and In-Context Learning
Alongside the theoretical accounts of the basis of ICL mentioned earlier, another line of research explains the origin of ICL through the distributional properties of the pre-training data. Chan et al. (2022) showed that ICL emerges when the training data exhibits particular distributional properties (such as burstiness, in which items appear in clusters rather than being uniformly distributed over time). Hahn and Goyal (2023) argued that ICL emerges from the compositional structures found in the pre-training data under the standard next-token prediction objective. Chen et al. (2024) found that parallel structures in the pre-training data give rise to ICL capability in LLMs. They defined parallel structures as pairs of phrases following similar templates in the same context window and found that removing parallel structures in the pre-training data significantly reduces LLMs’ ICL accuracy. Chen et al. also pointed out that despite the fact that the pre-training data is not formatted strictly as in-context prompts, i.e., input-output pairs, the naturalistic data often contains phrases following similar templates. Those phrase pairs could be conceptualized as in-context examples of implicitly defined, less structured shared “tasks”, such as copying n-grams, repeating syntactic constructions, and producing related types of world knowledge.
As structural priming is a well-attested phenomenon in humans, it is reasonable for us to hypothesize that structural priming is a factor that shapes the distribution of LLM pre-training data since humans tend to produce abundant parallel structures in naturalistic settings. For this reason, we view the repeated structures from structural priming as a case of the parallel structures in Chen et al.’s (2024) sense. Inspired by this data-centric perspective, we do not take ICL to necessarily involve explicit demonstration-answer pairs for specific tasks, as is typically understood in the literature. Instead, we conceptualize ICL as a more generalized notion that involves a sensitivity to parallelism through generic next-token prediction: any text in the context window will influence the model’s logits toward preferring similar text. This generalized notion of ICL, namely, the contextual influence of prompts in the context window, is analogous to priming in humans, as we elaborate in Section 3.1.
@@ 2.3 Structural Priming in Neural Language Models
Structural priming has been proposed as a means of probing the abstract mental representations of structural information in humans that define the notion of similarity underlying priming (Branigan and Pickering, 2017). Previous works have adopted this paradigm for probing learned linguistic representations in neural networks. van Schijndel and Linzen (2018) and Prasad et al. (2019) show that LSTMs (Gulordava et al., 2018) are capable of adapting to syntactic structures under fine-tuning. Specifically, these researchers fine-tune model weights on prime sentences and test target sentence probabilities in the updated model. This is a direct analog of the implicit learning account of structural priming as it involves weight updates. Recently, Sinclair et al. (2022) have shown that the GPT2 family Radford et al. (2019) shows robust structural priming without the use of explicit weight updates; when a prime sentence is concatenated to its corresponding target sentence, we see an increase in the probability the model assigns to the target sentence as compared to without a prime. [^1] Sinclair et al. (2022) have also demonstrated that the GPT2 models showed the lexical boost effect, another well-attested sub-phenomenon of structural priming, which is not our main focus here. Other works have demonstrated crosslingual structural priming in large language models Michaelov et al. (2023), suggesting that structural priming is robustly detected in LLMs. Jumelet et al. (2024) study in considerable details the linguistic factors the give rise to structural priming in current state-of-the-art LLMs.
In sum, there is a large body of work demonstrating that LLMs, like humans, show structural priming. This sets the stage for our current study of investigating the processing mechanisms underlying priming.
@@ 3.1 Overview of Our Approach
We first clarify our conceptualization of ICL. As stated in Section 2.2, instead of adopting the notion that ICL requires demonstration-answer pairs of a task as prompts in the context window, here we propose that any text in the context window will predispose the model toward producing similar text in the future. This text could take the form of demonstration-answer pairs, as is the case in the best-known examples of ICL, but it does not need to: it could be any type of text whose structure is then preferred by the model in subsequent predictions. Under this generalized framing, ICL can produce structural priming: a model that observes a sentence with a particular syntactic structure (say, the DO structure) can use ICL to replicate that syntactic structure, even though the sentence cannot be cleanly viewed as a demonstration-answer pair. That is, the less structured, implicitly defined “task” encoded by the DO sentence used as the prompt could be interpreted as “producing another sentence following the DO structural template exemplified in the prompt.”
Our research question is whether a gradient component is implicitly computed during the forward pass of processing the prompt in the generalized ICL setting. We investigate this question by testing whether LLMs show the IFE in the ICL setting. [^2] In work that was done concurrently with the present paper, Jumelet et al. (2024) provide evidence that LLMs show the IFE, corroborating this finding of our paper. The two papers differ in several ways. First is the main motivation: we use the IFE as a diagnostic for characterizing the type of mechanism that underlies ICL, while Jumelet et al. study the IFE for purposes of understanding how similar LLM structural priming is to human structural priming. A second difference stems from the analyses that are run: the current paper defines an explicit metric for the IFE in Section 3.6 based on each prime verb’s individual verb bias, while Jumelet et al. use a linear mixed model that involves a different metric of the verb biases (i.e., structural preference) to investigate the predictive power of various factors on the LLMs’ priming behaviors. Finally, the two papers differ in the form of evidence that they use for detecting the IFE: we show direct evidence for the IFE in LLMs through experimental design that isolates the effect of frequency, while Jumelet et al. present indirect evidence for the IFE in LLMs by analyzing which predictors in a linear mixed model have a significant effect: they show that the sentence-level surprisal and the structural preference are predictive of the priming effects, and they treat these two factors as evidence for the IFE in LLMs. As is illustrated in Figure 1, given that (i) it has been argued by psycholinguists that only error-driven learning mechanisms can give rise to the IFE; (ii) processing the prime sentence in the context window conditions the probability of the target sentence in the generalized notion of ICL; and (iii) standard structural priming in the ICL setting has been robustly observed , we hypothesize that the strength of the IFE will positively correlate with the strength of the ICL capability of LLMs: the stronger the ICL capability is, the better the gradient will be computed in the forward pass, leading to a stronger IFE.
@@
Specifically, we simulate structural priming across LLMs of various sizes with the two modes mentioned in Section 2.3. As is illustrated in Figure 2, the Fine-Tuning mode fine-tunes the LLM’s parameters on a single prime sentence, and the updated model is used to infer the probability of the target sentence. The Concatenation mode instead uses (our generalized notion of) ICL, where the prime sentence is directly concatenated with the target sentence, using the prime sentence as the prompt in the context window, and the probability of the target sentence is measured. The Fine-Tuning mode serves as a sanity check that LLMs are able to show the IFE when there is explicit error-driven, gradient-based learning. It sets the stage for our main focus: using the Concatenation mode to diagnose whether ICL involves implicit gradient-based learning.

### v2

# Is In-Context Learning a Type of Error-Driven Learning? Evidence from the Inverse Frequency Effect in Structural Priming

@@ Abstract
Large language models (LLMs) have shown the emergent capability of in-context learning (ICL). One line of research has claimed that ICL is functionally equivalent to gradient descent, a type of error-driven learning mechanism. In this paper, we introduce a new way of diagnosing whether ICL is functionally performing error-driven learning. Our approach is based on the inverse frequency effect (IFE)—a phenomenon in which an agent’s behavior is influenced to a greater degree when presented with improbable examples as compared to more likely ones. The IFE has previously been identified in psycholinguistics where humans exhibit the IFE in the context of structural priming (the tendency for people to produce sentence structures they have encountered recently). In that context, the IFE has been used as evidence that human structural priming must involve error-driven learning mechanisms. In our experiments, we simulated structural priming with ICL and found that LLMs indeed display the IFE, with the effect being stronger in larger models. We conclude that at least in the case we studied, ICL is indeed a type of error-driven learning, supporting the hypothesis that an error signal is implicitly computed in the forward pass during ICL. Our results suggest that both humans and LLMs make use of error-driven processing mechanisms in on-line processing. [^1] Code is available at: https://github.com/herbert-zhou/ICL_IFE
Is In-Context Learning a Type of Error-Driven Learning? Evidence from the Inverse Frequency Effect in Structural Priming
@@ 1 Introduction
To what extent do humans and language models use similar processing mechanisms? In some ways, language processing by language models and human learners appears to be substantially different. Humans display the ability to quickly and flexibly adapt to new contexts, while language models have historically required massive amounts of training data and a large number of parameters to exhibit human-like performance. Yet recent pre-trained large language models (LLMs) have shown the capacity to perform in-context learning (ICL): they flexibly adapt to novel tasks with only a small number of demonstrations provided as prompts in the context window without any parameter updates (Brown et al., 2020). This emergent capability could provide a way to bridge the divide between language models and humans: because ICL enables LLMs—like humans—to flexibly adapt to novel contexts, perhaps ICL shares important properties with the processing mechanisms used by humans.
@@
Within the body of research about why ICL arises and how it operates, one line of work has aimed to deepen our theoretical understanding of ICL by offering functional interpretations of ICL as a kind of implicit gradient descent during inference. Von Oswald et al. (2023) demonstrate that Transformer models, with appropriate choices of parameters, can process in-context demonstrations in a way that is functionally equivalent to performing gradient updates on the same demonstration examples. Garg et al. (2022), Zhang et al. (2024), and Ahn et al. (2024) show that standard Transformers (Vaswani et al., 2017) can be trained to implement learning algorithms for linear regressions under ICL-based training objectives. Dai et al. (2023) provide a mathematical construction showing a dual form between Transformer attention and gradient descent and interpreted ICL as a meta-optimization process that performs implicit fine-tuning. However, Shen et al. (2023) argue that the importance of these theoretical demonstrations is limited in that they do not take ICL to be an emergent property, instead assuming a training objective that optimizes for ICL. These demonstrations therefore deviate from actual LLMs pre-trained with natural data with a language modeling training objective. Indeed, Shen et al. find inconsistencies between ICL and gradient descent in real models, and therefore leave the equivalence between ICL and gradient descent as an open question.
@@
In this paper, we aim to better characterize what kind of learning mechanism ICL is by drawing a connection between ICL and human learning mechanisms, using a case study that allows us to evaluate off-the-shelf LLMs using natural language data. Specifically, we investigate whether ICL is a type of error-driven learning such that an error signal is implicitly computed during the forward pass. By focusing on this question, we take some first steps toward understanding whether ICL involves implicit gradient descent because gradient descent is one type of error-driven learning. To approach this question, we treat ICL as a processing mechanism of LLMs and borrow insights from methods of studying processing mechanisms in humans: we examine to what extent LLMs show the inverse frequency effect (IFE), a phenomenon in human structural priming (Branigan and Pickering, 2017) that has been argued to require an error-driven processing mechanism in humans, implicit learning (e.g., Chang et al., 2006). We focus our attention on the widely studied linguistic phenomenon of the dative alternation where the IFE has been robustly attested in humans. We demonstrate that LLMs not only show a robust IFE under standard fine-tuning (which involves explicit gradient descent, a type of error-driven learning), but also varying degrees of the IFE under the ICL setting, with larger models showing a stronger IFE. Given the correlation between ICL ability and model size, we use our results to conclude that, at least in the cases we study, ICL is indeed an error-driven learning mechanism.
@@
Overall, our results suggest that error-driven learning is an aspect of processing shared by humans and ICL-using LLMs. Our study has implications for both NLP/ML (1, 1, & 1) and linguistically-motivated analysis of LLMs (1 & 1):
@@
. We find evidence that ICL in off-the-shelf LLMs can be viewed as a form of error-driven learning in at least some cases.
[…]
. We show how the IFE can serve as a diagnostic for error-driven learning, paving the way for future work that can use this diagnostic to further investigate the circumstances under which ICL functions like error-driven learning.
@@
. We show that LLMs qualitatively display an important property of human language processing: the IFE in structural priming.
@@ 2 Background and Related Work
In this section, we lay out the building blocks necessary for motivating why we use the IFE to diagnose the error-driven nature of ICL. Our experimental approach is formally described in Section 3.1.
@@ 2.1 Structural Priming in Psycholinguistics
Encountering a syntactic structure can predispose speakers to repeat that structure in the near future, a phenomenon known as syntactic priming Bock (1986). For example, after encountering a double object (DO) sentence (e.g., Alice gave Bob a book), speakers tend to produce another DO structure (e.g., The student sent the professor a letter) rather than a semantically equivalent prepositional dative (PD) structure (e.g., The student sent a letter to the professor). Structural priming has been interpreted as an adaptation mechanism, where speakers adapt lexical and syntactic predictions to the current context Jaeger and Snider (2013), similar to the way in which LLMs adapt their outputs on the basis of prompts. Examples of priming typically involve two sentences; we will refer to them as the prime sentence and the target sentence, where the prime sentence comes first and influences the form of the target sentence.
@@
One important aspect of structural priming is the inverse frequency effect Jaeger and Snider (2008); Bernolet and Hartsuiker (2010); Kaschak et al. (2011): less-preferred syntactic alternatives cause stronger overall priming than more-preferred structures, where the degree to which a structure is preferred is operationalized as its relative frequency in the speaker’s experience. This can be seen by looking at sentences involving verbs with gradient structural preferences, or verb biases (or alternation biases; see Hawkins et al., 2020 for a systematic investigation of verb biases in neural models). The verb give allows both prepositional dative (PD) (The doctor gave a book to the judge) and double object (DO) (The doctor gave the judge a book) structures, but is biased toward (i.e., occurs more often with) DO in English. Under the IFE, a “prime” sentence involving give in its (less preferred) PD structure will cause a greater priming effect than a DO sentence, i.e., a PD prime will increase the probability of a PD occurring more than a DO prime will increase the probability of a DO occurring. The strength of PD priming (i.e., the increase in the probability of a PD target given a PD prime) inversely correlates with the expectation of a PD prime, determined by its verb bias Bernolet and Hartsuiker (2010).
@@
Two theories have been proposed to account for structural priming. Transient activation theory Pickering and Branigan (1998) claims that the activation of structural representations from the prime persists for a short time (in working memory) so that it is easier to reactivate the same structure at the next relevant opportunity. This form of transient activation theory does not however account for the IFE because the amount of residual activation is independent from verb biases and does not involve any error-driven mechanism. In contrast, implicit learning theory Chang et al. (2006) claims that humans are continuously and dynamically adjusting their probabilistic knowledge concerning the occurrence of grammatical structures (including verb biases) on the basis of their experience and use such information to predict the form of linguistic input. Crucially, under standard theories of learning, the update performed by the learner is error-driven, such that a larger update is performed in situations where the learner’s predictions are farther from the truth. In the context of priming, this would mean that priming strength is determined by the difference between the learner’s predictions and the actual prime sentence: the less the learner expects the observed prime structure, the larger the resulting error signal is that updates their expectations for that structure in the future, yielding a larger priming effect. Therefore, implicit learning, unlike transient activation, predicts the IFE. The two theories are not mutually exclusive and can co-exist to account for priming, a framing known as the dual mechanism account Tooley and Traxler (2010).
@@
In this study, we assume the correctness of the arguments from psycholinguistics that error-driven learning mechanisms are necessary to explain the IFE. Therefore, by examining whether LLMs show the IFE in the ICL setting, we can infer whether there is an implicit error signal computed in ICL.
Figure 1: An overview of our experimental design.
@@ 2.2 Structural Priming in Neural Language Models
Structural priming has been used as a tool for probing the mental representations of structure in humans, under the assumption that these representations define the notion of similarity underlying priming (Branigan and Pickering, 2017). Previous works have adopted this paradigm for probing learned linguistic representations in neural networks. van Schijndel and Linzen (2018) and Prasad et al. (2019) show that LSTMs (Gulordava et al., 2018) are capable of adapting to syntactic structures under fine-tuning. Specifically, these researchers fine-tune model weights on prime sentences and test target sentence probabilities in the updated model. This is a direct analog of the implicit learning account of structural priming as it involves weight updates. Recently, Sinclair et al. (2022) have shown that the GPT2 family Radford et al. (2019) shows robust structural priming without the use of explicit weight updates; when a prime sentence is concatenated to its corresponding target sentence, the model assigns a higher probability to the target sentence than it does when the target sentence appears without a prime. Other works have demonstrated crosslingual structural priming in large language models Michaelov et al. (2023), suggesting that structural priming is robustly detected in LLMs. In concurrent work, Jumelet et al. (2024) provide a detailed study of the linguistic factors that give rise to structural priming in current state-of-the-art LLMs, including presenting evidence that LLMs display the IFE; our work differs from theirs in that we study the IFE as a diagnostic for error-driven learning, while Jumelet et al.’s goal is to compare LLM priming with human priming. In sum, there is a large body of work demonstrating that LLMs, like humans, show structural priming. This sets the stage for our current study of investigating the processing mechanisms underlying priming.
@@ 3.1 Overview of Our Approach
We first clarify our conceptualization of ICL. In most of the literature, ICL is conceptualized as involving explicit demonstration-answer pairs. However, Chen et al. (2024) find that parallel structures in pre-training data substantially contribute to ICL capabilities in LLMs, where parallel structures are defined as pairs of phrases that appear in the same context window and that follow similar templates. Although these parallel structures are not explicitly framed as input-output pairs for a task, they can be conceptualized as in-context examples of less-structured implicit “tasks” such as copying n-grams, repeating syntactic constructions, and producing related types of world knowledge. Inspired by this perspective, we conceptualize ICL as a more generalized notion that involves a sensitivity to parallelism through generic next-token prediction: any text in the context window will influence the model’s logits toward preferring similar text. This generalized notion of ICL, namely, the contextual influence of prompts in the context window, is analogous to priming in humans. The fact that Chen et al. (2024) showed that parallel structures in pre-training substantially contribute to the emergence of ICL supports the view that we can frame ICL in these more general, parallelism-focused terms. Under this generalized framing, ICL can produce structural priming: a model that observes a sentence with a particular syntactic structure (say, the DO structure) can use ICL to replicate that syntactic structure, even though the sentence cannot be cleanly viewed as a demonstration-answer pair. That is, the less structured, implicitly defined “task” encoded by the DO sentence used as the prompt could be interpreted as “producing another sentence following the DO structural template exemplified in the prompt.”
Our research question is whether an error signal is implicitly computed during the forward pass of processing the prompt in the generalized ICL setting. We investigate this question by testing whether LLMs show the IFE in the ICL setting. Given that (i) it has been argued by psycholinguists that only error-driven learning mechanisms can give rise to the IFE; (ii) processing the prime sentence in the context window conditions the probability of the target sentence in our generalized notion of ICL; and (iii) standard structural priming in the ICL setting has been robustly observed , we hypothesize that the strength of the IFE will positively correlate with the strength of the ICL capability of LLMs: the stronger the ICL capability is, the better the error signal will be computed in the forward pass, leading to a stronger IFE. In particular, given recent empirical evidence showing that larger models have stronger ICL capabilities (e.g., Wei et al., 2023; Dong et al., 2024; Chen et al., 2025), we further hypothesize that the scaling of both model size and data size contributes to LLMs’ ICL capability. Thus, we test the hypothesis that ICL becomes more like error-driven learning with both model scale and data scale.
@@
We simulate structural priming across LLMs of various sizes with the two previously explored methods mentioned in Section 2.2 (see Figure 1). The Fine-Tuning mode updates the LLM’s parameters on the basis of a single prime sentence, and the updated model is used to infer the probability of the target sentence. The Concatenation mode instead uses (our generalized notion of) ICL, where the prime sentence is concatenated with the target sentence, so that the prime constitutes the prompt/context at the point in which the probability of the target sentence is measured. The Fine-Tuning mode serves as a sanity check that LLMs are able to show the IFE with explicit error-driven learning. Because of the nature of gradient descent, changes to model weights during fine-tuning are a function of the degree to which a prime sentence is correctly predicted. A more unexpected prime will therefore give rise to larger gradients and updates, as well as a larger priming effect. Once we have demonstrated this result empirically, the stage will then be set for our main focus: using the Concatenation mode to diagnose whether ICL involves implicit error-driven learning.